# Tensor tree learns hidden relational structures in data to construct generative models

### v1

@@ Abstract
Based on the tensor tree network with the Born machine framework, we propose a general method for constructing a generative model by expressing the target distribution function as the quantum wave function amplitude represented by a tensor tree. The key idea is dynamically optimizing the tree structure that minimizes the bond mutual information. The proposed method offers enhanced performance and uncovers hidden relational structures in the target data. We illustrate potential practical applications with four examples: (i) random patterns, (ii) QMNIST hand-written digits, (iii) Bayesian networks, and (iv) the stock price fluctuation pattern in S&P500. In (i) and (ii), strongly correlated variables were concentrated near the center of the network; in (iii), the causality pattern was identified; and, in (iv), a structure corresponding to the eleven sectors emerged.
tensor tree, generative modeling, network structure optimization, mutual information
Generative models thrive on the adaptability of architectures tailored to the data’s characteristics. The architecture is often chosen manually, such as using RNNs for time series and sequential data. However, in some cases, the network structure is automatically selected. For example, a fully connected neural network is optimized by pruning parameters, such as the weight matrices, to reflect the correlations among data elements [1]. Typically, the architecture of a generative model is expressed as a network that represents connections between data elements. A high performance is achieved by manually or automatically optimizing this network structure to suit the data characteristics. Therefore, learning the relational structure of the target data is important to obtain better generative models.
Recently, a method of using the quantum state measurement process as a generative model was proposed [2]. From Born’s rule, it follows that the probability of obtaining a measurement output x is the squared amplitude of the wave function of a quantum state. This probability model for generative modeling is called the Born machine. In [2] and [3], a tensor network was proposed to represent the wave function in the Born machine. A tensor network is illustrated as a network of tensors, with a connection representing the partial contraction between tensors. It has been utilized in various machine learning methods, such as classifiers [4], neural networks [5], and reinforcement learning [6]. In [2] and [3], a one-dimensional chain (called tensor train) and a balanced tree were adopted for the tensor network representing a wave function, respectively. The differences in the performance of resulting generative models suggest the importance of choosing the network structure. However, how we can choose the best network structure for a given data remains unclear.
@@
Results
Optimization of tensors and network structure in a Born machine with a tensor tree
@@ Results
where $|\psi_{\theta}\rangle$ is the quantum state of the machine. Here, we replace a real quantum system with a tensor network with tree structure to represent $\langle\textbf{x}|\psi_{\theta}\rangle$. Generally, a tensor network is a partial contraction of the product of tensors. It is depicted as a graph, where the nodes (i.e., circles or ellipses) represent tensors and the lines their indices. A line connecting two nodes indicates an index to be contracted, referred to as a bond. The dimension of an index, i.e., how many possible values it may take, is called the bond dimension. We consider a tensor network with no loop, which we call tensor tree. Each component tensor has three indices.
In the Born machine with a tensor tree, the mutual information that measures mutual dependence between two groups of outputs has a strict limit. By cutting a bond in a tensor tree, we decompose it into two subtrees. In addition, outputs are divided into two groups, one for each subtree. For example, for the network in Fig.1(b), by cutting a thick bond, the machine’s outputs are divided into parts A and B. We call the mutual information between A and B as bond mutual information (BMI) and is given as $I(A,B)$. $I(A,B)$ is bounded from above by the entanglement entropy of $|\psi_{\theta}\rangle$ for the same bipartition[7, 8]. In addition, the entanglement entropy for the bipartition by a bond cut is less than the logarithm of the bond dimension of the cut bond[9]. Therefore, $I(A,B)$ is smaller than the logarithm of the bond dimension, $\chi$:
@@
.1 Examples
@@
.1.1 Random patterns and images of handwritten digits
@@
Generally, we would like a generative model to memorize finite random bit patterns accurately while effectively capturing the strong correlation among the bits. Although, generative modeling with a tensor train has struggled to learn long random sequences, in [3], it was demonstrated that these issues can be resolved by a balanced tensor tree, indicating the importance of the right choice of the tensor network in addressing this problem. We consider a particularly challenging problem with the target data representing a long-range correlation to explore this issue further. Here, the data to be learned consists of 10 sequences of 128 bits. While the left- and right-most 32 bits are generated as mutually independent random binary numbers, the bits in the intermediate part are fixed at 0 for all ten patterns, as illustrated in a tensor train in Fig.2(a). Since ten is less than the total number of possible bit patterns, it creates a strong correlation between the left- and right-most parts.
@@
Figure 2: Results of random binary patterns and images of handwritten digits. (a) Tensor train with open branches corresponding to random bit variables, represented by rainbow-colored circles. Bits with dotted open branches in a middle region are fixed at 0. (b) Negative log-likelihood in learning processes for ten random binary patterns. The initial network is a tensor train. (c) Optimized network structure for ten random binary patterns. Colored circles represent random bit variables. The color of an edge indicates the BMI value. (d) Optimized network structure for the images of the handwritten digits. The color of an edge indicates the amplitude of the bond mutual information. The initial network is a random tensor tree. (e) Converged values of negative log-likelihood for the images of the handwritten digits vs. bond dimension. (f) Ranking of pixels according to their distance from the center of the tensor tree network.
@@
Starting with the tensor train as the initial tensor network, we compare two cases: (i) the network fixed as the initial tensor train, and (ii) the network modified by the ATT method. When the network is adaptively modified, the negative log-likelihood (NLL) for random patterns converges to the expected lower bound, i.e., $\ln(10)$(Fig.2(b)). However, the NLL of tensor trains converges to local minimum values larger than $\ln(10)$. Fig.2 (c) shows a particular realization of the final network structure. The color of the circles represents their original positions, as shown in Fig.2(a). The red and blue circles, initially far apart, are tightly clustered in the final tree, reflecting their strong correlation. The irrelevant intermediate bits (blueish green or ocher) are excluded from it and barely connected (i.e., connected by low-BMI edges).
@@
In the previous example, the relevant random variables are separated in the tensor train, making the train structure disadvantageous compared to the structure derived by branch-reconnection. Below, we consider another scenario where the relational structure among the random variables is not immediately apparent, namely, 2D images. The balanced tensor tree outperforms the tensor train for images because its network structure aligns with the geometrical relationship among pixels [3]. However, to construct a chain or balanced tree of random bits (pixels), we need to know their spatial location, i.e., they require some prior knowledge about the target data. Here, we compare the present method with fixed network methods under the condition of no prior knowledge. We apply them to images of handwritten digits from the QMNIST dataset [10] using an initial random tensor tree or random pixel permutation (see the METHODS section for the details), which eliminates prior knowledge of the dataset’s implicit geometrical relationship.
@@
Fig.2(e) displays the final values of NLL for the test handwritten digit images. Comparing the results of the fixed balanced tree with (crosses) and without (solid circles) permutation shows the significant effect of the prior knowledge about the spatial arrangement of the pixels. In contrast, network optimization with branch-reconnection (blue and red) shows performance comparable to the balanced tree for the bond dimension around 10, even with the initial random tree. Fig.2(d) illustrates an optimized tensor tree network. The edges with strong BMI are concentrated near the center of the optimized network. We measure the distance from the center to an open edge representing a pixel to analyze the properties of the optimized network. In Fig.2(f), the color map shows pixel ranking based on the distance from the center. The compact region near the network’s center is also the central region in the 2D images. Similar to the case of the separated random sequences, the generative model automatically learns the relevant relational structure among the random variables and places them close together on the tensor network.
.1.2 Random variables of Bayesian networks
@@
We tested the ATT method using the data generated from three Bayesian polytree networks. In Fig.3(a-c), the Bayesian network is shown by nodes with digits and arrows. The optimized tensor tree is illustrated by circles and bonds. A node with a digit represents a random variable corresponding to an open index of the tensor tree. An arrow indicates the cause-to-effect direction between two random variables. The network in Fig.3(a) is a simple sequence with no branching. Fig.3(b) and (c) depict scenarios containing a branch and collision in the Bayesian network structure, respectively. The optimized tensor tree, starting from random tensor trees, successfully captures the corresponding topology of Bayesian networks for all cases, though our results do not detect the direction of dependencies explicitly.
.1.3 Stock price fluctuation patterns in S&P 500 index
@@
To evaluate the effectiveness of the ATT method for real data with an unknown structure, we built a generative model for the fluctuation pattern of stock prices. We focused on the fluctuation pattern of stock prices listed on S&P 500, a widely recognized stock price index. We analyzed 436 stocks that have been a part of S&P 500 index since 2010. We convert the change rate of each stock price into a binary value: 1 if it is higher than the average for all stocks and 0 otherwise. Thus, our dataset comprises 3589 samples, each represented by a 436-dimensional vector with binary components. To evaluate the generalization ability of the resulting generative models, we use half of the sample for training and the other half for evaluating the generalization ability by measuring the NLL.
@@
Although the stocks are categorized into 11 sectors and even finer classifications according to the “Global Industry Classification System,” we do not feed the generative model with such prior knowledge. We start from a random tree and let the tree evolve using the branch-reconnection algorithm described above. Fig.3(d) represents the NLL values after learning has converged with a given bond dimension. For comparison, we also conduct training without branch-reconnections (i.e., sticking to the initial random tree throughout learning). A clear improvement achieved by the branch-reconnections can be observed in the NLL for training and testing data sets. The discrepancy between the training and testing data is the measure of the generalization performance (the smaller the better). The figure shows that the generalization performance of the generative model is at its maximum around the bond dimension 4. The bond dimension larger than five shows over-fitting.
@@
Figure 3: (a-c) Target Bayesian networks of random binary variables and corresponding tensor network structures: (a) single dependency with no branching, (b) single dependency with branching, and (c) multiple dependencies. In each diagram, the upper layer is the target Bayesian network, and the lower layer is the corresponding tensor network. All tensor trees are not schematic diagrams but actual solutions obtained by the method regardless of the initial network configurations. (d,e) The result of adaptive tensor tree generative modeling applied to the stock price fluctuation patterns in S&P 500 index: (d) Bond-dimension dependency of the negative log-likelihood, and (e) sample of generated tree structure at the bond dimension of 5. Companies are colored according to the sector to which they belong.
Discussion

### v2

@@ Abstract
Based on the tensor tree network with the Born machine framework, we propose a general method for constructing a generative model by expressing the target distribution function as the amplitude of the quantum wave function represented by a tensor tree. The key idea is dynamically optimizing the tree structure that minimizes the bond mutual information. The proposed method offers enhanced performance and uncovers hidden relational structures in the target data. We illustrate potential practical applications with four examples: (i) random patterns, (ii) QMNIST handwritten digits, (iii) Bayesian networks, and (iv) the pattern of stock price fluctuation pattern in S&P500. In (i) and (ii), the strongly correlated variables were concentrated near the center of the network; in (iii), the causality pattern was identified; and in (iv), a structure corresponding to the eleven sectors emerged.
Keywords: generative modeling, Born machine, tensor network, tensor tree, network structure optimization, mutual information
1 Introduction
Generative models thrive on the adaptability of architectures tailored to the data’s characteristics. Learning the relational structure of the target data is important to obtain better generative models. The architecture of a generative model can be represented as a network that illustrates the connections between various data elements. High performance is achieved by optimizing this network structure to align with the characteristics of the data. However, the selection of the network structure is often done manually; for instance, recurrent neural networks are common for time series and sequential data. Pruning the parameters of weight matrices, as discussed in [1], is one approach for automatically selecting the network structure.
Recently, there has been extensive research on generative modeling on quantum computers. The Born machine[2] is a generative model that utilizes the projective measurement of the quantum state. Born’s rule dictates that the probability of obtaining a measurement output is the squared amplitude of the wave function of a quantum state. The theoretical quantum advantage offered by Born machines on parametrized quantum circuits was discussed in [3, 4, 5]. Born machines are promising applications of quantum computers[6, 7]. By utilizing a tensor network representation of a wave function[8, 9], we can efficiently build Born machines even on classical computers, which has been quite successful for various test datasets[2, 10].
A tensor network represents the partial contraction between tensors as a connection in a network. It has been utilized in various machine learning methods, such as classifiers [11], neural networks [12], and reinforcement learning [13]. In [2] and [10], a one-dimensional chain (called tensor train) and a balanced tree were adopted for the tensor network representing a wave function, respectively. The differences in the performance of resulting generative models suggest the importance of choosing the network structure. However, how we can choose the best network structure for a given data remains unclear.
@@ 1 Introduction
An optimized tensor network can also resolve significant optimization issues, the prevalence of barren plateaus in optimization landscapes, for the Born machine based on a quantum circuit[14]. Consequently, our findings improve the potential of quantum computing in practical applications.
2 Method
2.1 Optimization of tensors and network structure in a Born machine with a tensor tree
@@
where $|\psi_{\theta}\rangle$ is the quantum state of the machine. Here, we replace a real quantum system with a tensor network with a tree structure to represent $\langle\textbf{x}|\psi_{\theta}\rangle$. Generally, a tensor network is a partial contraction of the product of tensors. It is depicted as a graph, where the nodes (i.e., circles or ellipses) represent tensors and the lines their indices. A line connecting two nodes indicates an index to be contracted, referred to as a bond. The dimension of an index, i.e., how many possible values it may take, is called the bond dimension. We consider a tensor network with no loop, which we call tensor tree. Each component tensor has three indices. The tensor train and balanced tree used in [2] and [10], respectively, are special cases of the general tree structure. Based on a tree structure, we can efficiently calculate various exact contractions and marginal distributions using recursive procedures.
In the Born machine with a tensor tree, the mutual information that measures mutual dependence between two groups of outputs has a strict limit. By cutting a bond in a tensor tree, we decompose it into two subtrees. In addition, outputs are divided into two groups, one for each subtree. For example, for the network in Fig.1(b), by cutting a thick bond, the machine’s outputs are divided into parts A and B. We call the mutual information between A and B as the bond mutual information (BMI) and is given as $I(A,B)$. This quantity, $I(A,B)$ is bounded from above by the entanglement entropy of $|\psi_{\theta}\rangle$ for the same bipartition[15, 16]. In addition, the entanglement entropy for the bipartition by a bond cut is less than the logarithm of the bond dimension of the cut bond[8, 9]. Therefore, $I(A,B)$ is smaller than the logarithm of the bond dimension, $\chi$:
@@
2.2 Branch-reconnection algorithm
The details of our branch-reconnection in Fig.1(b) are as follows:
  1. 1.
    Initialize the network and all tensor elements.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
    Replace the 4-leg tensor with the chosen decomposition.
  6. 6.
    If the termination condition is satisfied, terminate.
  7. 7.
    Move to a bond connected to one of the two new tensors and return to step (3).
In steps (3) and (4), we improve the combined tensor and new tensors in the decomposition by employing gradient-descent updates to minimize the negative log-likelihood(NLL)[2, 10]. Assuming that the target distribution of generative modeling is an empirical data distribution, $\pi_{\textrm{emp}}$, NLL is defined as
where ${\cal M}$ is the ensemble of data samples and $D_{\textrm{KL}}$ is a Kullback-Leibler divergence. We perform a single gradient-descent update in step (3) and ten updates in step (4) of the procedure in Sec. 2.2. The learning rate is set to 0.001 in this study if we do not change.
The complexity of this algorithm is $O(\chi^{5})$, in which $\chi$ is the maximum bond dimension. It equals that of the two-site update of the fixed tree network method[10].
We also utilized a reconnection-based decomposition of a composite tensor to calculate the ground state of a quantum model with a tensor tree[17, 18, 19].
2.3 Canonical form of tree tensor network
We can define a canonical form for a tensor tree [20] to determine the optimal truncation from local calculations. We first choose the root edge to construct the canonical form of a given tensor tree. Then, all the tensors become isometries by recursively applying singular value decomposition (SVD) starting from the leaves. Finally, we obtain the matrix $\Lambda$ on the root edge. The canonical form of a tensor tree consists of a central weight on the root edge and isometries directed to a root.
After we optimize the composite tensor defined at a root edge to reduce an NLL for data, we can decompose it into a central weight and two isometries connected to the root edge with a restricted rank by the SVD. The decomposition by SVD is globally optimal because the other tensors are isometries in the canonical form.
Using SVD, the root’s position can be moved to any neighboring edge. Our method chooses the edge to be updated as the root. After modifying the local connection and updating the tensors around the root edge, we move the root position to the neighboring edge that has not been updated for the longest time.
The Born machine is practically advantageous because there is no need to restrict the parameters to ensure the positivity of the distribution function. This is because the squared amplitude of the wave function represents it. Hence, we can use SVD to precisely decompose the whole distribution function into a tensor tree with the canonical form. Interestingly, this might contribute to the stability of our optimization process because the SVD is more stable for decomposing a tensor than the non-negative matrix factorization[21].
2.4 Statistical estimation of mutual information
$I(A,B)=\sum_{(a,b)}p(a,b)\ln\frac{p(a,b)}{p(a)p(b)},$ (4)
where $a$ and $b$ are configurations on sub-systems $A$ and $B$, respectively, $p(a,b)=p_{\theta}(\textbf{x}=(a,b))$ is a joint distribution, and $p(a)=\sum_{b}p(a,b)$ and $p(b)=\sum_{a}p(a,b)$ are marginal distributions. We can efficiently compute the values of $p(a)$ and $p(b)$ utilizing a tensor tree representation of a joint distribution in the canonical form. Due to the orthogonality of isometries, we can replace the isometries with identity matrices on a path from a root to open indexes of marginalized variables for constructing a marginal distribution. However, estimating the mutual information in a tensor tree state is challenging due to the need to calculate the summation for all configurations $(a,b)$. The number of configurations grows exponentially with the system size.
where ${\cal M}^{\prime}$ is a set of data samples. The estimator (5) may be negative if the empirical data distribution $\pi_{\textrm{emp}}(a,b)$ differs from the joint distribution $p(a,b)$ in the early stage of optimization process. Then, we replace it with the absolute value. Such modification has no significant effect on the optimization process.
An alternative method involves estimating the average of $\ln[p(a,b)/p(a)p(b)]$ using Monte Carlo sampling. If a Born machine has a tree structure, it can efficiently calculate the marginal distribution of a random variable given the other random variable states. The conditional distribution is then used to sample random variables perfectly. Repeating these steps can generate an independent sample $(a,b)$ with probability $p(a,b)$. Consequently, we can efficiently do Monte Carlo estimation of (4) as (5) with ${\cal M}^{\prime}$, the ensemble of samples generated from $p(a,b)$ directly.
Due to the low quality of the Born machine in its early stages, we employed the first method outlined in (5) to estimate mutual information. The standard deviation of the estimation of BMI with empirical distribution is almost within $2\sim 3$% of the average of the BMI.
In order to reduce the computational cost, we used the same set of data samples to estimate NLL and mutual information. Therefore, when the number of training data samples is large, we utilized the same mini-batch, a randomly selected subset of data samples.
3 Results
@@
3.1 Random patterns and images of handwritten digits
@@
Generally, we would like a generative model to memorize finite random bit patterns accurately while effectively capturing the strong correlation among the bits. Although generative modeling with a tensor train has struggled to learn long random sequences, in [10], it was demonstrated that these issues can be resolved by a balanced tensor tree, indicating the importance of the right choice of the tensor network in addressing this problem. We consider a particularly challenging problem with the target data representing a long-range correlation to explore this issue further. Here, the data to be learned includes ten sequences, each 128 bits long. The set of the ten sequences serves as a training batch. While the left- and right-most 32 bits are generated as mutually independent random binary numbers, the bits in the intermediate part are fixed at 0 for all ten patterns, as illustrated in a tensor train in Fig.2(a). Since ten is less than the total number of possible bit patterns, it creates a strong correlation between the left- and right-most parts. In this case, we change the learning rate in the gradient-descent update in step (3) and (4) of the procedure in Sec. 2.2 to 0.05 and set the number of iterations for termination at step (6) of the procedure to 3000.
@@
Figure 2: Results of random binary patterns and images of handwritten digits. (a) Tensor train with open branches corresponding to random bit variables, represented by rainbow-colored circles. Bits with dotted open branches in a middle region are fixed at 0. (b) Negative log-likelihood in learning processes for ten random binary patterns. The initial network is a tensor train. (c) Optimized network structure for ten random binary patterns. Colored circles represent random bit variables. The color of an edge indicates the amplitude of the bond mutual information. (d) Converged values of negative log-likelihood for the images of the handwritten digits vs. bond dimension. (e) Optimized network structure for the images of the handwritten digits. The color of an edge indicates the amplitude of the bond mutual information. The initial network is a random tensor tree. (f) Ranking of pixels according to their distance from the center of the tensor tree network.
@@
Starting with the tensor train as the initial tensor network, we compare two cases: (i) the network fixed as the initial tensor train, and (ii) the network modified by the ATT method. When the network is adaptively modified, NLL for random patterns converges to the expected lower bound, i.e., $\ln(10)$(Fig.2(b)). However, the NLL of tensor trains converges to local minimum values larger than $\ln(10)$. Fig.2 (c) shows a particular realization of the final network structure. The color of the circles represents their original positions, as shown in Fig.2(a). The red and blue circles, initially far apart, are tightly clustered in the final tree, reflecting their strong correlation. The irrelevant intermediate bits (blueish green or ocher) are excluded from it and barely connected (i.e., connected by low-BMI edges).
@@
In this example, the relevant random variables are separated in the tensor train, making the train structure disadvantageous compared to the structure derived by branch-reconnection. Below, we consider another example where the relational structure among the random variables is not immediately apparent, namely, 2D images. The balanced tensor tree outperforms the tensor train for images because its network structure aligns with the geometrical relationship among pixels [10]. However, to construct a chain or balanced tree of random bits (pixels), we need to know their spatial location, i.e., they require some prior knowledge about the target data. Here, we compare the present method with fixed network methods under the condition of no prior knowledge. We apply them to images of handwritten digits from the QMNIST dataset [22] using an initial random tensor tree or random pixel permutation, which eliminates prior knowledge of the dataset’s implicit geometrical relationship.
The QMNIST dataset is an improved version of the Mixed National Institute of Standards and Technology database (MNIST) dataset[23], a collection of handwriting digits from 0 to 9 on $28\times 28$ pixels. In QMNIST, the number of test images increases to the number of learning images, 60000. The intensity of a pixel ranges from 0 to 255. As in [10], by padding zero intensity pixels around an original image, we embed an original image into $32\times 32$ pixels. We make a binary bit from a pixel’s intensity. If the intensity is larger than 127, the bit is set to one; otherwise, it is set to zero. The final image consists of $32^{2}$ bits. For data samples ${\cal M}$ to estimate NLL and mutual information, we use a mini-batch of size 1000. However, if the mini-batch size is small, instability occurs in estimating mutual information. We set the interval for changing the mini-batch to 1000 gradient-descent updates. The number of iterations is set to $10^{6}$ until the pseudo-program in Sec. 2.2 is terminated at step (6).
We generated an initial random tree in this study as follows:
  1. 1.
  2. 2.
  3. 3.
    Until the list S contains only one element, repeat step (2).
The final element in S is a tree with the random variables associated with its leaves. We estimated the variance of NLLs from the five initial random trees.
@@
Fig.2(d) displays the final values of NLL for the test handwritten digit images. Comparing the results of the fixed balanced tree with (crosses) and without (solid circles) permutation shows the significant effect of the prior knowledge about the spatial arrangement of the pixels. In contrast, network optimization with branch-reconnection (blue and red) shows performance comparable to the balanced tree for the bond dimension around 10, even with the initial random tree. Fig.2(e) illustrates an optimized tensor tree network. The edges with strong BMI are concentrated near the center of the optimized network. We measure the distance from the center to an open edge representing a pixel to analyze the properties of the optimized network. In Fig.2(f), the color map shows pixel ranking based on the distance from the center. The compact region near the network’s center is also the central region in the 2D images. Similar to the case of the separated random sequences, the generative model automatically learns the relevant relational structure among the random variables and places them close together on the tensor network.
As shown in Fig.2(b) and (d), the ATT’s results show much lower NLL than the fixed tensor network models for both cases. This means that when we cannot use prior knowledge, such as the two-dimensional arrangement of the pixels, the structure optimization is efficient in reducing the NLL.
3.2 Random variables of Bayesian networks
@@
We tested the ATT method using the data generated from three Bayesian polytree networks. In Fig.3(a-c), the Bayesian network is shown by nodes with digits and arrows. A node with a digit represents a random variable corresponding to an open index of the tensor tree. An arrow indicates the cause-to-effect direction between two random variables. The network in Fig.3(a) is a simple sequence with no branching. Fig.3(b) and (c) depict scenarios containing a branch and collision in the Bayesian network structure, respectively. In these Bayesian networks, each random variable is binary. We generate the target data according to the causal relation. When there is a single cause, say $A$, we generate the result bit $B$ with the conditional probability $P(B=A|A)=r$. In contrast, when there are multiple causes, we first calculate the intermediate bit using the exclusive OR(XOR) of all the cause bits and generate the result bit with the same conditional probability. The present article’s correlation rate $r$ is fixed to be 0.8. For data samples, we used a mini-batch of size 1000 and updated it when every edge had been examined once with the previous mini-batch. In all cases, we use a bond dimension of 4 for a tensor tree. The optimized tensor tree is illustrated by circles without digits and bonds in Fig.3(a-c). Since the NLL for test data fluctuates in the optimization process when changing a mini-batch, we showed the optimized tensor tree with the lowest NLL for test data in Fig.3(a-c). We did $3000$ iterations until the pseudo-program in Sec. 2.2 is terminated at step (6). The optimized tensor tree, starting from random tensor trees, successfully captures the corresponding topology of Bayesian networks for all cases, though our results do not detect the direction of dependencies explicitly.
We note that the ATT method finds a many-body correlation among variables “7,” “15,” and “16” in Fig.3(c) because the variable “16” stochastically depends on the XOR of variables “7” and “15.” Because the XOR cannot be represented as a two-point correlation, most conventional methods based only on the two-point correlations would fail to obtain the correct topology for Fig.3(c).
In the context of decoherence in a quantum circuit, the structure of a Bayesian network for measurement results mirrors that of the quantum circuit itself[28, 29]. In these examples, the ATT method effectively infers the underlying structure from the data.
3.3 Stock price fluctuation patterns in S&P 500 index
@@
To evaluate the effectiveness of the ATT method for real data with an unknown structure, we built a generative model for the fluctuation pattern of stock prices. We focused on the fluctuation pattern of stock prices listed on S&P 500, a widely recognized stock price index. We analyzed 436 stocks that have been a part of S&P 500 index since 2010. We convert the change rate of each stock price into a binary value: 1 if it is higher than the average for all stocks and 0 otherwise. Thus, our whole dataset comprises 3589 samples, each represented by a 436-dimensional vector with binary components. To evaluate the generalization ability of the resulting generative models, we randomly select half of the sample as a training batch and the other half to evaluate the generalization ability by measuring the NLL.
@@
Although the stocks are categorized into 11 sectors and even finer classifications according to the “Global Industry Classification System,” we do not feed the generative model with such prior knowledge. We start from a random tree and let the tree evolve using the branch-reconnection algorithm in Sec. 2.2 with $2\times 10^{6}$ iterations until it is terminated at step (6). Fig.3(d) represents the NLL values after learning has converged with a given bond dimension. For comparison, we also conduct training without branch-reconnections (i.e., sticking to the initial random tree throughout learning). A clear improvement achieved by the branch-reconnections can be observed in the NLL for training and testing data sets. The discrepancy between the training and testing data is the measure of the generalization performance (the smaller the better). The figure shows that the generalization performance of the generative model is at its maximum around the bond dimension 4. The bond dimension larger than five shows over-fitting.
@@
Figure 3: (a-c) Target Bayesian networks of random binary variables and corresponding tensor network structures: (a) single dependency with no branching, (b) single dependency with branching, and (c) multiple dependencies. In each diagram, the lower layer is the target Bayesian network, and the upper layer is the corresponding tensor network. All tensor trees are not schematic diagrams but actual solutions obtained by the method regardless of the initial network configurations. (d,e) The result of the adaptive tensor tree method applied to the stock price fluctuation patterns in S&P 500 index: (d) Bond-dimension dependency of the negative log-likelihood, and (e) sample of generated tree structure at the bond dimension of 5. Companies are colored according to the sector to which they belong.
4 Discussion and Conclusion